How to find the slope (or rate of change) between the the two points with coordinates (1997, 1.26) and (2006, 2.31)
Answer by stanbon(75887)

How to find the slope (or rate of change) between the the two points with coordinates (1997, 1.26) and (2006, 2.31)
------------------------------------
slope = (2.31-1.26)/(2006-1997) = 1.05/9 = 0.1167
Cheers,
Stan H.